I need help on this problem. I need to know where to put parentheses to make this equation true.
16/4+2*9/3 = 6
It seems the only way to solve is guess and check, but I could not find a solution. Please help me?

Found 2 solutions by rfer, MathTherapy:
Answer by rfer(10417)

16/4+2*9/3=6
(16/4)+((2*9)/3)=6

Answer by MathTherapy(639)

16/4+2*9/3 = 6

I can't find any either. Are you sure that the "+" between the "4" and "2" isn't a division sign? Check the problem again, and resubmit if you erroneously stated this problem.

David